OpenStudy (blackcat):

A square has a side of 6.25 feet. What is the area in square feet? Round to the correct significant digits.

Directrix (directrix):

A = L * W where length and width are the same A = 6.25 * 6.25 A = ? @blackcat

OpenStudy (blackcat):

?

Directrix (directrix):

What is the product of 6.25 and 6.25 ?